MICROWAVE SPECTROSCOPY OF BARIUM 5d_{3/2}nℓ AUTOIONIZING STATES

\underline{EDWARD SHUMAN} and TOM GALLAGHER, Department of Physics, University of Virginia, Charlottesville, VA 22903.

We have observed microwave resonance transitions between the autoionizing Ba 5d_{3/2}ng J = 2 and 5d_{3/2}(n + 1)h J = 3 states for 45 ≤ n ≤ 49. These measurements, made between 50 and 75 GHz, lead to quantum defects of +0.019(2) and +0.003(2) for this pair of ng and nh states. These measurements demonstrate that the core polarization model, widely used to characterize non penetrating Rydberg states, fails to describe Rydberg states converging to anisotropic or non s cores. We believe that this is the first time spectroscopy of autoionizing states using microwaves has been performed.